MIXING MATHEMATICS AND MATERIALS

DAVID HOFFMAN

Abstract. Recent uses of differential geometry in materials science are reviewed here, in particular the September issue of the Phil. Trans. Royal Soc., entitled “Curvature and chemical Structure.”